# FAID: Fine-Grained AI-Generated Text Detection Using Multi-Task Auxiliary and Multi-Level Contrastive Learning

### v2

@@ 4.3 Multi-Task Auxiliary Learning
Multi-task learning Caruana (1997) allows a model to learn several tasks concurrently by sharing relevant information across them. This joint learning process helps the model develop more general and distinctive features. Therefore, it improves the model’s generalizability to new data. Building on the previously described contrastive learning framework, we extend the encoder by attaching an MLP classifier at its output layer. This classifier is responsible for binary classification, determining whether a given text was written by a human or a LLM. Let the probability of $i^{th}$ sample with label $x_{i}=0$ be $p_{i}$. To train this component, we apply a cross-entropy loss function, denoted as $\mathcal{L}_{ce}$, and defined as follows:
@@ 4.4 Handling Unseen Data without Retraining
Unseen data, whether from an unseen domain or an unfamiliar generator, remains a significant challenge even for state-of-the-art LLM-generated text detection methods, due to the increasing capabilities of LLMs. We tried to use the model classifier on its own, but we ended up using a vector database along with Fuzzy k-Nearest Neighbors, as illustrated in Figure 2. The results were written in Appendix F. Specifically, when dealing with unseen data, we use our model to embed these samples and add them to our existing vector database. Through careful parameter tuning, this approach enables our system to effectively handle newly encountered unseen data without the need retraining.
@@ 5 Experiments
Scenario Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
In-domain,
Known generator
@@
Unseen domain,
Unseen generator
@@
Unseen
generators
LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
@@
Unseen domain
and Unseen
generators
LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
@@
Table 1: Performance with three labels. For in-domain and known generator datasets, we train the detectors on the training set and evaluate them on the test set. For unseen data evaluation, we use the detector trained on FAIDSet and evaluate on the unseen datasets. The best results are in bold and the second best are underlined.
5.1 Datasets
LLM-DetectAIve Abassy et al. (2024) encompasses various domains, including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews. The original human-written and machine-generated texts were augmented by multiple LLMs to create a 236k-example dataset with two new labels: (imachine-written then machine-humanized, and (iihuman-written then machine-polished.
HART Bao et al. (2025) with 21,500 examples including student essays, arXiv abstracts, story writing, and news articles. It contains four categories: human-written, AI-refined, AI-generated, and humanized AI-generated texts. The authors further expanded the dataset by creating additional instances with unbalanced label distributions.
Unseen domain: We created a dataset consisting of 150 IELTS Writing essays from Kaggle ibrahimmazlum (2023), where all original texts are human-written. We used these essays to generate human–LLM collaborative and LLM-generated texts with the same models with FAIDSet.
@@ 5.1 Datasets
FAIDSet LLM-DetectAIve 0.7596 0.7697 0.7690 0.7653
T5-Sentinel 0.7568 0.7985 0.7840 0.7837
SeqXGPT 0.6941 0.6802 0.6420 0.6603
FAID 0.7964 0.8328 0.8352 0.8327
LLM-DetectAIve LLM-DetectAIve 0.9049 0.9064 0.8352 0.8693
T5-Sentinel 0.8154 0.8137 0.8009 0.8105
SeqXGPT 0.8712 0.7890 0.7608 0.7441
FAID 0.9089 0.8817 0.8672 0.8737
HART LLM-DetectAIve 0.8900 0.8787 0.8674 0.8715
T5-Sentinel 0.7852 0.7713 0.7834 0.7759
SeqXGPT 0.6470 0.5582 0.4540 0.5075
FAID 0.8996 0.9157 0.8648 0.8667
@@ 5.2 Baselines
We adapted the method of Abassy et al. (2024) by fine-tuning a roberta-base sequence classification model. The input texts were tokenized using the RoBERTa tokenizer with a maximum sequence length of 256, and the model was trained for three-class detection.
@@
We adapt the T5-Sentinel framework introduced by Chen et al. (2023) for our three-class setup. Following the original configuration, we train using the AdamW optimizer (batch size 128, learning rate $1\times 10^{-4}$, weight decay $5\times 10^{-5}$). This allows direct comparison with prior T5-based detectors under our experimental conditions.
@@
Following Wang et al. (2023) by modeling token-level likelihood patterns from LLM tokenizers for sentence-level detection, we adapted their method by updating the tokenizer models to align with our dataset’s label space. This adjustment ensures that the extracted token-level log-probability features correspond more accurately to the model types represented in our data.
5.3 Human, LLM or Human-LLM?
Table 1 shows the performance of FAID and three baselines in three evaluation settings.
FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domains, (iii) unseen generators, and (iv) unseen domain & generators settings, followed by LLM-DetectAIve in (i) and (iv), and T5-Sentinel in (ii) and (iii). Although SeqXGPT aimed at extracting sequence-level features, it struggles with texts from advanced models, whose coherent and human-like writing styles reduce detectable distinctions.
FAID improves generalization performance over unseen domains and generators compared to others. We can see that generalizing to (ii) unseen domains and (iv) unseen domain & generator remains challenging, with accuracy of 62.78% and 66.55%, respectively. These results suggest that FAID is an effective method for addressing the multilingual fine-grained LLM-generated text detection task, improving performance by leveraging multi-level contrastive learning to capture generalizable stylistic differences tied to LLM families, rather than overfitting to surface-level artifacts.
@@ 5.4 Identifying Different Generators
We further examine the detector to identify specific AI generators seen during training. The goal of FAID is not only to detect AI involvement, but also to identify the specific LLM or family applied, treating them as distinct authors. As shown in Table 2, across the three datasets, FAID consistently achieves higher performance compared to other baselines in almost all metrics. LLM-DetectAIve achieves comparable results with our detector FAID on the test set of LLM-DetectAIve dataset, except that its precision is slightly lower. FAID’s high performance when dealing with text originating from diverse known generators within these datasets indicates that FAID learned unique writing patterns and features of different AI generators by leveraging multi-level contrastive learning.
@@ 6 Conclusion and Future Work
We presented FAIDSet, a multilingual fine-grained LLM-generated text detection dataset comprising 84k examples, and FAID, a framework designed to distinguish between human-written, LLM-generated, and human–LLM collaborative texts. FAID integrates multi-level contrastive learning and multi-task auxiliary objectives, treating LLM families as stylistic “authors,” which enables it to capture subtle linguistic and stylistic cues that generalize effectively across domains and evolving generative systems. Moreover, its Fuzzy KNN-based inference and training-free incremental adaptation contribute to strong robustness and adaptability to unseen data.
@@
In the future, we plan to extend FAIDSet to cover more languages, generators, and domains, particularly informal genres such as social media and student writing, to enhance cross-lingual and domain robustness further. We also aim to incorporate adversarial LLM-generated texts and more complex forms of human–LLM collaboration to capture the evolving dynamics of AI-assisted text creation.
7 Limitations
While FAID demonstrates strong performance and generalization across various domains and LLMs, several limitations remain. First, although our dataset is multilingual and multi-domain, it is still limited in terms of low-resource languages and niche writing domains, which may affect performance in those contexts. FAIDSet is synthetic by construction. The controlled generation enables causal-style analysis, but it under-represents the messy, tool-chain-specific edits, wich are seen in the wild. We mitigate this with diverse prompt paraphrases, manual spot-checks, and generalization tests to held-out generators. Second, our framework is based on the observation that texts produced by LLMs from the same family share similar stylistic features. However, this may break down when a single text is influenced by multiple LLM systems – such as when a human uses different models for drafting, rewriting, and polishing a text. In such cases, the resulting style may blend traits from multiple LLMs, making it more difficult to attribute authorship to a single LLM family or clearly distinguish collaboration boundaries.
8 Ethics and Broader Impact
@@
A primary ethical consideration is the data license. We reused existing datasets for our research: LLM-DetectAIve, HART, and IELTS Writing, which have been publicly released under clear licensing agreements. We adhere to the intended usage of these dataset licenses.
@@
FAIDSet streamlines both the creation and the rigorous testing of FAID. By spotting LLM-generated material, FAID helps preserve academic integrity, flag potential misconduct, and protect the genuine contributions of authors. More broadly, it supports efforts to prevent the misuse of generative technologies in areas such as credential falsification. Detecting LLM-generated content across different languages can be tricky, due to language’s grammar and style. By enabling robust, multilingual and multi-generator detection with accurate results, FAIDSet empowers people everywhere, especially in academic scenarios, to deploy AI responsibly. At the same time, it fosters critical digital literacy, giving everyone a clear understanding of both the strengths and limits of generative AI.
@@
FAID is designed to enhance transparency in AI-assisted writing by enabling the fine-grained detection of AI involvement in text generation. While this has clear benefits for academic integrity, and content provenance, we acknowledge the potential for misuse. For instance, such tools could be used to unfairly penalize individuals in educational or professional settings based on incorrect or biased predictions. To mitigate this, we stress that FAID is not intended for high-stakes decision-making without human oversight.
@@ Appendix A Some Current Datasets on AI-generated Texts
We analyze many current datasets on AI-generated text and summarize them in Table 3. Many datasets contain fine-grained labels and multi-generators, but all of the analyzed datasets are monolingual – only in English. This problem prompts us to collect and formulate a new dataset for multilingual, multi-domain, and multilingual AI-generated text, enabling us to adapt the approach to real-world scenarios.
@@
Appendix B FAIDSet Statistics and Analysis
B.1 Statistics
Our FAIDSet includes 83,350 examples, which are divided into three subsets: train, validation, and test, with the ratios shown in Table 4. The data set also comprises various sources of human-written texts, which are described in Table 5.
Subset Human AI human–LLM collab.
@@ B.1 Statistics
Source No. Human texts
@@
HUST theses (Vietnamese) 11,164
@@ B.2 Diverse Prompt Strategies
To avoid biasing our generated corpora toward any single style or topic, we employ a broad set of prompt templates when synthesizing LLM-generated and human–LLM collaborative texts. By varying prompt structures, content domain, and complexity, we ensure that the resulting outputs cover a wide distribution of writing patterns, vocabulary, and rhetorical devices. This diversity helps our detector generalize more effectively to real-world inputs.
Concretely, we use five prompts for LLM-generated texts in Table 6 and several categories of human–LLM collaborative texts in Table 7, Table 8 and Table 9 consecutively:
@@
Table 6: List of diverse prompt templates used to generate FAIDSet – Label: LLM-generated
@@
Table 7: List of diverse prompt templates used to generate FAIDSet – Label: LLM-polished
@@
Table 8: List of diverse prompt templates used to generate FAIDSet – Label: LLM-continued
@@
Table 9: List of diverse prompt templates used to generate FAIDSet – Label: LLM-paraphrased

### v3

@@ 4.3 Multi-Task Auxiliary Learning
Multi-task learning Caruana (1997) allows a model to learn several tasks concurrently by sharing relevant information across them. This joint learning process helps the model develop more general and distinctive features. Therefore, it improves the model’s generalizability to new data. Building on the previously described contrastive learning framework, we extend the encoder by attaching an MLP classifier at its output layer.
This classifier performs binary classification, determining whether a given text was written by a human or an LLM. Let the probability of $i^{th}$ sample with label $x_{i}=0$ be $p_{i}$. To train this component, we apply a cross-entropy loss function, denoted as $\mathcal{L}_{ce}$, and defined as follows:
@@ 4.4 Handling Unseen Data without Retraining
Unseen data, whether from an unseen domain or an unfamiliar generator, remains a significant challenge even for state-of-the-art LLM-generated text detection methods, as LLMs continue to improve. We tried to use the model classifier on its own, but we ended up using a vector database along with Fuzzy k-Nearest Neighbors, as illustrated in Figure 2. The results are given in Appendix F. Specifically, when dealing with unseen data, we use our model to embed these texts and add them to our existing vector database. Through careful parameter tuning, this approach enables our system to effectively handle newly encountered unseen data without retraining.
@@ 5 Experiments
Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
@@
Table 1: Performance with three labels. The best results are in bold and the second best are underlined.
5.1 Datasets
LLM-DetectAIve Abassy et al. (2024) encompasses various domains, including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews. We augmented the original labels human-written and machine-generated using multiple LLMs to create a 485,405-example dataset with two new labels: (imachine-written then machine-humanized, and (iihuman-written then machine-polished.
HART Bao et al. (2025) has 21,500 examples, including student essays, arXiv abstracts, story writing, and news articles. It covers four categories: human-written, AI-refined, AI-generated, and humanized AI-generated texts. The authors further expanded the dataset by creating additional instances with unbalanced label distributions.
Unseen domain: We created a dataset consisting of 150 IELTS essays from Kaggle [^3] https://www.kaggle.com/datasets/mazlumi/ielts-writing-scored-essays-dataset, where all texts are human-written. We used these essays to generate human–LLM collaborative and LLM-generated texts with the same models with FAIDSet.
Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
@@ 5.1 Datasets
Unseen generators LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
@@
Unseen domains and Unseen generators LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
@@
Table 2: Performance with three labels with unseen data. We use the detector trained on FAIDSet and evaluate on the unseen datasets. The best results are in bold and the second best are underlined.
@@
FAIDSet LLM-DetectAIve 75.96 76.97 76.90 76.53
T5-Sentinel 75.68 79.85 78.40 78.37
SeqXGPT 69.41 68.02 64.20 66.03
FAID 79.64 83.28 83.52 83.27
LLM-DetectAIve LLM-DetectAIve 90.49 90.64 83.52 86.93
T5-Sentinel 81.54 81.37 80.09 81.05
SeqXGPT 87.12 78.90 76.08 74.41
FAID 90.89 88.17 86.72 87.37
HART LLM-DetectAIve 89.00 87.87 86.74 87.15
T5-Sentinel 78.52 77.13 78.34 77.59
SeqXGPT 64.70 55.82 45.40 50.75
FAID 89.96 91.57 86.48 86.67
@@ 5.2 Baselines
We adapted the method of Abassy et al. (2024) by fine-tuning a roberta-base sequence classification model. We tokenized the input texts using the RoBERTa tokenizer with a maximum sequence length of 256, and we trained the model for three-class detection.
@@
We adapted the T5-Sentinel framework introduced by Chen et al. (2023) for our three-class setup. Following the original configuration, we trained using the AdamW optimizer (batch size 128, learning rate $1\times 10^{-4}$, weight decay $5\times 10^{-5}$). This allows direct comparison with prior T5-based detectors under our experimental conditions.
@@
We adopted the method of Wang et al. (2023), who model token-level likelihood patterns from LLM tokenizers for sentence-level detection: we updated the tokenizer models to align with our dataset’s label space. This adjustment ensures that the extracted token-level log-probability features better reflect the model types in our data.
5.3 Human-Only, LLM-Only, or Human-LLM?
Tables 1 and 2 show the performance of FAID and three baselines in three evaluation settings.
As shown in Table 1, FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domains, (iii) unseen generators, and (iv) unseen domain & generators settings. It is followed by LLM-DetectAIve for (i) and (iv), and by T5-Sentinel for (ii) and (iii). Despite being designed to extract sequence-level features, SeqXGPT struggles with texts from advanced models, whose coherent, human-like writing styles reduce the detectable distinctions.
FAID further improves generalization performance over unseen domains and generators compared to others as illustrated in Table 2. We can see that generalizing to (ii) unseen domains and (iv) unseen domain & generator remains challenging, with accuracy of 62.78% and 66.55%, respectively. These results suggest that FAID is an effective method for addressing the multilingual fine-grained LLM-generated text detection task, improving performance by leveraging multi-level contrastive learning to capture generalizable stylistic differences tied to LLM families, rather than overfitting to surface-level artifacts.
@@ 5.4 Identifying Different Generators
The goal of FAID is not only to detect whether AI was used to produce the target text, but also to identify the specific LLM family, treating the families as distinct authors. As shown in Table 3, FAID consistently achieves higher performance compared to other baselines in almost all evaluation measures.
LLM-DetectAIve achieves results comparable to FAID on the test set of its dataset, except that its precision is slightly lower. FAID’s high performance when dealing with text from diverse known generators in these datasets indicates that it learned unique writing patterns and features of different generators by leveraging multi-level contrastive learning.
Language Accuracy Precision Recall F1-macro
English 96.41 96.02 95.59 95.77
Vietnamese 94.42 95.60 94.22 94.76
Table 4: Language-wise performance on FAIDSet.
5.5 Language-wise Performance
Table 4 presents the classification performance of FAID on each language of FAIDSet. The results show that FAID achieves consistently strong performance across both subsets. While performance on English is slightly higher, FAID maintains competitive accuracy and F1-score on Vietnamese, indicating that the model does not rely solely on high-resource language features.
5.6 Generalizability to Unseen Human–LLM type of Collaboration
To assess FAID’s robustness to unseen collaborative writing patterns between humans and LLMs, we conducted an additional experiment focusing on hybrid authorship styles that were not present in the original training dataset.
We conducted manual quality control to enhance the reliability of our dataset and to ensure the robustness of the FAID model. A team of five annotators, all IT-majored, fluent in English, and aged 18–25, participated in the annotation process. Each annotator was assigned approximately 80 samples, covering all collaborative writing styles, for a total of 400 human-reviewed instances across various collaboration modes.
During this manual revision stage, the annotators followed two key quality control principles:
  1. 1.
    Ensuring logical and informational consistency between the outputs and the original texts, where the output length was controlled to be 70–150% of the source length.
  2. 2.
    Improving quality through spelling correction, synonym replacement, and word refinement to ensure natural fluency and stylistic coherence.
FAID was evaluated on this manually curated dataset without further fine-tuning. Despite these samples representing an unseen distribution, the model achieved a strong overall accuracy of 84.8%, precision of 82.8%, and recall of 85.0% across all collaboration categories. This suggests a strong generalization capability beyond the data distributions encountered during model development. It also highlights FAID’s sensitivity to fine-grained stylistic blending, where human revision only partially conceals the generative footprint of LLMs.
5.7 Generalizability to Real-World Scenarios
To further assess FAID’s generalizability beyond the controlled benchmark setting, we conducted an additional user study simulating realistic academic and professional writing scenarios. The goal was to evaluate whether FAID can maintain detection accuracy when applied to authentic, unconstrained text produced through real interactions with LLMs.
A group of five volunteers with diverse academic backgrounds was instructed to engage with four popular AI systems: ChatGPT, Gemini, DeepSeek, and Llama 3.1 to generate text resembling authentic student or professional writing.
The interactions were open-ended but were guided by three types of collaboration commonly encountered in two real-world use cases: writing a paper summary (as an abstract) and writing a passage for their own graduation thesis.Each participant was asked to generate five outputs per model, yielding a total of 200 real-world text samples. To ensure authenticity and diversity, the participants were encouraged to adjust the prompts iteratively, to include follow-up clarifications, and to perform light editing, mimicking realistic human–LLM co-writing behavior.
FAID achieved strong performance, with overall accuracy of 88.5%, precision of 85.9%, and recall of 89.7%, indicating strong generalizability to unseen real-world scenarios despite being trained only in an in-domain setting.
@@ 6 Conclusion and Future Work
We presented FAIDSet, a multi-domain, multilingual fine-grained LLM-generated text detection dataset comprising 83k examples, and FAID, a framework designed to distinguish between human-written, LLM-generated, and especially, human–LLM collaborative texts in practice.
FAID integrates multi-level contrastive learning and multi-task auxiliary objectives, treating LLM families as stylistic “authors”, which enables it to capture subtle linguistic and stylistic cues that generalize effectively across domains and evolving generative systems. Moreover, its Fuzzy k-Nearest Neighbors-based inference and training-free incremental adaptation contribute to strong robustness and adaptability to unseen data.
@@
In future work, we plan to extend FAIDSet to cover more languages, generators, and domains, particularly low-resource languages and informal genres such as social media and student writing, to further enhance cross-lingual and domain robustness. We further plan to incorporate adversarially LLM-generated texts as well as more complex forms of human–LLM collaboration in order to better capture the evolving dynamics of AI-assisted text creation.
Limitations
While FAID demonstrates strong performance and generalization across various domains and LLMs, several limitations remain. First, although our dataset is multilingual and multi-domain, it remains limited in low-resource languages and niche writing domains, which may affect performance in those contexts. FAIDSet is synthetic by construction. The controlled generation enables causal-style analysis, but it under-represents the messy, tool-chain-specific edits seen in the wild. We mitigate this with diverse prompt paraphrases, manual spot-checks, and generalization tests to held-out generators. Second, our framework is based on the observation that texts produced by LLMs from the same family share similar stylistic features. However, this may break down when a single text is influenced by multiple LLMs, e.g.,  when a human uses different models for drafting, rewriting, and polishing. In such cases, the resulting style may blend traits from multiple LLMs, making it more difficult to attribute authorship to a single LLM family or clearly distinguish collaboration boundaries.
Ethics and Broader Impact
@@
A primary ethical consideration is the data license. We reused existing datasets for our research: LLM-DetectAIve, HART, and IELTS Writing, which have been publicly released with clear licenses and well-documented terms of use. We adhere to the intended usage of these datasets.
@@
FAIDSet streamlines both the creation and the rigorous testing of FAID. By spotting LLM-generated material, FAID helps preserve academic integrity, flag potential misconduct, and protect the genuine contributions of authors. More broadly, it supports efforts to prevent the misuse of generative technologies, such as credential falsification. Detecting LLM-generated content across different languages can be tricky, due to the language’s grammar and style. By enabling robust, multilingual, and multi-generator detection with accurate results, FAIDSet empowers people everywhere, especially in academic scenarios, to deploy AI responsibly. At the same time, it fosters critical digital literacy, giving everyone a clear understanding of both the strengths and the limits of generative AI.
@@
FAID is designed to enhance transparency in AI-assisted writing by enabling the fine-grained detection of AI involvement in text generation. While this has clear benefits for academic integrity and content provenance, we acknowledge the potential for misuse. For instance, such tools could be used to unfairly penalize individuals in educational or professional settings based on incorrect or biased predictions. To mitigate this, we stress that FAID is not intended for high-stakes decision-making without human oversight.
@@ Appendix A Some Current Datasets on AI-generated Texts
We review existing datasets for AI-generated text detection and summarize their key characteristics in Table 5. While many of these datasets provide fine-grained labels and cover multiple text generators, all are monolingual and limited to English. This limitation highlights a critical gap in current resources and motivates the construction of a new multilingual, multi-domain, and multi-generator dataset for AI-generated text detection, which is essential for developing methods that generalize to real-world, cross-lingual scenarios.
@@
Subset Human LLM Human–LLM
collab.
@@
Source Human Texts
@@
HUST theses (Vietnamese) 11,159
@@
Appendix B FAIDSet Statistics and Analysis
B.1 Statistics
Our FAIDSet includes 83,350 examples, which are divided into three subsets: train, validation, and test, with the ratios shown in Table 6. The dataset also comprises various sources of human-written text, as described in Table 7.
@@ B.2 Diverse Prompt Strategies
In order to avoid biasing our generated corpora toward some style or topic, we use a broad set of prompt templates when synthesizing LLM-generated and human–LLM collaborative texts. By varying prompt structures, content domains, and complexity, we ensure that the resulting outputs cover a wide variety of writing patterns, vocabulary, and rhetorical devices. This diversity helps our detector generalize more effectively to real-world inputs.
Concretely, we use five prompts for LLM-generated texts in Table 8 and several categories of human–LLM collaborative texts in Tables 9, 10, and 11, consecutively:
@@
Table 8: List of diverse prompt templates used to generate FAIDSet – Label: LLM-generated.
@@
Table 9: List of diverse prompt templates used to generate FAIDSet – Label: LLM-polished.
@@
Table 10: List of diverse prompt templates used to generate FAIDSet – Label: LLM-continued.
@@
Table 11: List of diverse prompt templates used to generate FAIDSet – Label: LLM-paraphrased.